OpenStudy (anonymous):

What are the solutions of x2 + 2x − 24 = 0? x = 4, x = −6 x = −4, x = −6 x = 4, x = 6 x = −4, x = 6

OpenStudy (anonymous):

It is the last one correct @hartnn ?

hartnn (hartnn):

yup :)

OpenStudy (anonymous):

Thanks

hartnn (hartnn):

no wait

hartnn (hartnn):

its not

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

@Luis_Rivera That's what I got ...

hartnn (hartnn):

 so its (x-4)(x+6)=0 x-4=0 and x+6=0 x=4 and x=-6